OpenStudy (anonymous):

use the sum/difference identities to evaluate exactly cos 150 degree

OpenStudy (anonymous):

\[\cos 150^0 = \cos(90^0 + 60^0)\] \[= \cos(90^0 + 60^0) = \cos90^0 \cos 60^0 - \sin90^0 \sin60^0\] \[= 0 \times \frac{1}{2} - 1 \times \frac{\sqrt 3}{2} = 0 - \frac{\sqrt 3}{2}\] \[= - \frac{\sqrt 3}{2}\] Hence; \[\cos 150^0 =- \frac{\sqrt 3}{2}\] Formula used cos(A+B) = cos A cos B- sin A sin B @shorouq91

OpenStudy (anonymous):

@shorouq91

OpenStudy (anonymous):

why i cant use sin(45-30)

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

Sure but you wiil have to convert cos 150 into sin @shorouq91

OpenStudy (anonymous):

cos105 degree

OpenStudy (anonymous):

\[\cos150^0=\cos(180^0-30^0)= -\cos30^0 = - \frac {\sqrt3}{2}\] @shorouq91

OpenStudy (anonymous):

Another way:

OpenStudy (anonymous):

the question is cos 105 degree

OpenStudy (anonymous):

\[\cos150^0=\cos(90^0+60^0)= -\sin60^0 = - \frac {\sqrt3}{2}\] @shorouq91

OpenStudy (anonymous):

@shorouq91

OpenStudy (anonymous):

i got the wrong question